### v6

@@ 4 Simulation evidence
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with the best feasible bias correction (gold squares), QR with the baseline feasible bias correction (gold crosses), and QR with infeasible bias correction (gold dashed line) for DGP1 with $n\in\{50,100,200\}$. The error bands (gold bars) correspond to $\Phi^{-1}(1-0.05/12)\times$ MCSE to account for the joint testing of 6 hypotheses. (Notice that the scaled MCSE of the estimator grows with $n$.) All results are based on 40,000 simulation repetitions.
Figure 3: Bias (multiplied by $n$) before and after correction for DGP2–DGP4
@@
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with the best feasible bias correction (gold squares), QR with the baseline feasible bias correction (gold crosses), and QR with infeasible bias correction (gold dashed line) for DGP2–DGP4 with $n=100$. We use classical QR for DGP2 and DGP3 and exact IVQR (implemented via the MILP formulation in Appendix D) for DGP4. The error bands (gold bars) correspond to $\Phi^{-1}(1-0.05/12)\times$ MCSE to account for the joint testing of 6 hypotheses. The DGP2 and DGP3 results are based on 40,000 simulation repetitions; the DGP4 results are based on 10,000 repetitions. For DGP4, the infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
@@
Notes: Panel (a) compares the RSME for classical QR without (blue) and with (gold) bias correction. Panel (b) compares the RMSE for exact IVQR (implemented via the MILP formulation in Appendix D) without (blue) and with (gold) bias correction. The results are based on 40,000 simulation repetitions.
@@ 5 Empirical application
The second-order bias matters most in applications with small sample sizes. We therefore illustrate our bias correction approach using the classical dataset of Engel (1857), analyzed by Koenker and Bassett (1982) and Koenker and Hallock (2001), among others. The data contain information on annual income and food expenditure (in Belgian francs) for $n=235$ Belgian working-class households and are obtained from the R package quantreg (Koenker, 2022). One feature of these data is the growing dispersion of the outcome variable (food expenditure) as a function of the regressor (income) (Koenker and Hallock, 2001), which is similar to our Monte Carlo designs. We divide the values of income and food expenditure by $1000$ so that the unit of measurement becomes a thousand Belgian francs. This makes the scale of the variables comparable to that in our Monte Carlo simulations, allowing us to use the same baseline bandwidth choices.
We estimate classical QRs of food expenditure ($Y$) on income ($W$) and a constant (blue dots). The bias-corrected estimators (gold squares) are obtained using the baseline bandwidth choice $h_{1,n}=A_{G}n^{-1/5}$, $h_{2,n}=A_{Q}n^{-1/7}$, and $h_{3,n}=A_{\kappa}n^{-2/15}$ with $(A_{G},A_{Q},A_{\kappa})=(1,1,1)$.
Figure 5 presents the results. Panel (a) shows the impact of bias correction. The results suggest that bias correction is more important at the upper tail, where we find differences between the classical QR and the bias-corrected estimates. After the bias correction, we observe more heterogeneity across quantile levels, suggesting that the second-order bias leads classical QR to underestimate this heterogeneity. Notice that there is a second-order bias even at the median (i.e., least absolute deviation) regression.
To investigate the sensitivity of our results to the choice of the three bandwidths, we also report the minimum and maximum bias-corrected estimates when $(A_{G},A_{Q},A_{\kappa})$ varies over the set $\mathcal{A}$ in equation (10). Our results suggest that bias-corrected estimates are insensitive to the bandwidth choice.
The main takeaway from the bias decomposition is that while all three components play a role, the sample moments term (i) can be very large and accounts for most of the bias at the upper tail quantiles. The QR estimates with the highest overall second-order bias have the largest contribution from the sample moments term ($\tau\in\{0.5,0.7\}$). These results suggest that one could reduce the worst-case bias substantially by only correcting the sample moment term, which is akin to a 1-step Newton correction (see Appendix E).
@@
Notes: Panel (a) compares the classical QR estimates with (blue dots) and without (gold squares) bias correction, where $(A_{G},A_{Q},A_{\kappa})=(1,1,1)$. The bars indicate the minimum and the maximum of the estimates when varying $(A_{G},A_{Q},A_{\kappa})$ over the grid $\mathcal{A}$ defined in (10). Panel (b) shows the contributions of different bias components to the overall second-order bias.
@@ 6 Conclusion
We demonstrate that QR and IVQR estimators can exhibit a non-negligible second-order bias. We theoretically characterize this bias and use our results to derive a novel feasible bias correction method. Our method can effectively reduce the second-order bias at a very low computational cost and without substantially increasing RMSE.
Acknowledgements
@@ A.2 Auxiliary results for exact estimators
[…]
@@
[…]
@@ B.1 Auxiliary results
We use the maximal inequality in Lemma 1 of Ota et al. (2019) to establish (26). To do so, we verify the three conditions of this lemma. First,
@@
[…]
@@ B.2 Proofs of main results on bias correction
$\hat{\kappa}_{\tau}$ $=\kappa_{\tau}+O_{p}\left(\frac{1}{\sqrt{n}}\right)+O_{p}\left(\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right)+o_{p}\left(\frac{\log h_{n}}{n^{5/6}h_{n}}\right)+O(h_{n}^{2})+O\left(\frac{\|\hat{G}-{G}\|}{h_{n}}\right).$
Since, under the optimal step size for the estimator $\hat{G}$, we have $O(\|\hat{G}-{G}\|/h_{n})=O_{p}\left(h^{-1}_{n}n^{-2/5}\sqrt{\log n}\right)$, the overall rate is
$\max\left\{n^{-1/2},\,\frac{\sqrt{\log n}}{n^{2/5}h_{n}},\,\frac{\log h_{n}}{n^{5/6}h_{n}},\,h_{n}^{2},\,\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right\}.$
The optimal bandwidth is $h_{3,n}\propto n^{-2/15}$ with
$\hat{\kappa}_{\tau}=\kappa_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{4/15}}\right).$
@@ Appendix F Additional figures
Figure 6: Bias (multiplied by $n$) before and after correction for DGP1, sensitivity to bandwidth choice
(a) $A_{G}\pm 0.5$
(b) $A_{Q}\pm 0.5$
(c) $A_{\kappa}\pm 0.5$
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with feasible bias correction using the baseline bandwidth choice $A=(1,1,1)$ (gold crosses), QR with feasible bias correction using ±$\pm 0.5$ deviation in one bandwidth at a time (up-pointing triangles for $+0.5$, down-pointing triangles for $-0.5$), and QR with infeasible bias correction (gold dashed line) for DGP1. In Panel (a), we vary $A_{G}$, in Panel (b), we vary $A_{Q}$, and in Panel (c), we vary $A_{\kappa}$. All results are based on 40,000 simulation repetitions.
@@
Notes: Panel (a) compares the MAD for classical QR without (blue) and with (gold) bias correction. Panel (b) compares the MAD for exact IVQR (implemented via the MILP formulation in Appendix D) without (blue) and with (gold) bias correction. The results for both DGPs are based on 40,000 simulation repetitions.
Figure 8: Confidence interval coverage before and after bias correction for DGP1, different sample sizes
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope for classical QR without bias correction (blue dots) and QR with the best feasible bias correction (gold crosses) for DGP1 with $n\in\{50,100,200\}$. All results are based on 40,000 simulation repetitions.
Figure 9: Confidence interval coverage before and after correction for DGP2–DGP4
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope for classical QR without bias correction (blue dots) and QR with the best feasible bias correction (gold crosses) for DGP2–DGP4 with $n=100$. All results are based on 40,000 simulation repetitions.

### v7

@@ 4.3 Results
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and QR with infeasible bias correction (gold dashed line) for DGP1–DGP3. All results are based on 5,000 simulation repetitions.
Figure 3: Bias (multiplied by $n$) before and after correction for DGP4–DGP6
@@
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for IVQR (implemented via the MILP formulation in Appendix D) without bias correction (blue dots), IVQR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and IVQR with infeasible bias correction (gold dashed line) for DGP4–DGP6. All results are based on 5,000 simulation repetitions. The infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
@@
Notes: The panels compare the RMSE for estimators without bias correction (blue), with infeasible bias correction (grey) and with feasible bias correction based on the rule-of-thumb bandwidth choice (gold) for (a) DGP1, classical QR and (b) DGP4, IVQR. All results are based on 5,000 simulation repetitions.
Figure 5: Confidence interval coverage before and after correction for DGP1 and DGP4
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP1 (classical QR) and DGP4 (IVQR). All results are based on 5,000 simulation repetitions.
@@ 5 Empirical application
The second-order bias matters most in applications with small sample sizes. We therefore illustrate our bias correction approach using the classical dataset of Engel (1857), analyzed by Koenker and Bassett (1982) and Koenker and Hallock (2001), among others. The data contain information on annual income and food expenditure (in Belgian francs) for $n=235$ Belgian working-class households and are obtained from the R package quantreg (Koenker, 2022). One feature of these data is the growing dispersion of the outcome variable (food expenditure) as a function of the regressor (income) (Koenker and Hallock, 2001), which is similar to our Monte Carlo designs. We divide the values of income and food expenditure by $1000$ so that the unit of measurement becomes a thousand Belgian francs. This makes the scale of intercept and slope parameters comparable.
@@
Notes: Panel (a) compares the classical QR estimates (blue dots) to the bias-corrected QR estimates (gold squares) with 90% confidence intervals (bars). The bias correction is based on the rule-of-thumb bandwidth choice with $(\tilde{A}_{G},\tilde{A}_{Q},\tilde{A}_{\kappa})=(2,1.5,2)$. Panel (b) shows the contributions of the different bias components to the overall second-order bias.
We estimate classical QRs of food expenditure ($Y$) on income ($W$) and a constant (blue dots). The bias-corrected estimates (gold squares) are obtained using the recommended rule-of-thumb bandwidth choice th $(\tilde{A}_{G},\tilde{A}_{Q},\tilde{A}_{\kappa})=(2,1.5,2)$. Figure 6 presents the results. Panel (a) compares the classical and the bias-corrected QR estimates. The results suggest that the impact of bias correction is more pronounced around the median and in the tails. The magnitude of the differences between the original and bias-corrected estimates can be larger than 50% of the standard errors, which is substantial given that we are focusing on the second-order bias.
The main takeaway from the bias decomposition is that while all three components play a role, the sample moment term (i) and especially the Hessian term (iii) can be large and account for most of the bias. The $\kappa$-term (ii) is smaller overall and only matters in the tails.
@@ 6 Conclusion
We demonstrate that classical QR and IVQR estimators can exhibit a non-negligible second-order bias. We characterize this bias theoretically and use this characterization to derive a novel analytical bias correction method. The proposed feasible bias correction reduces the bias of QR and IVQR estimators across a variety of settings at a very low computational cost. However, there is scope for further improving its performance when the sample size is very small and the instruments are weak. The simulation performance of the infeasible bias correction based on the population version of our theoretical bias formula suggests that exploring alternative feasible bias correction approaches is a promising direction for future research. For example, one could consider regularization approaches or explore imposing parametric assumptions to improve the estimation of the bias components.
Acknowledgments
@@ A.2 Auxiliary results for exact estimators
Under Assumptions 1.1, 2 and 3, $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\theta_{\tau}\|=o_{p}\left(1\right)$, where $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ for any $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,\text{QR}}$.
@@
Under Assumptions 13, for any exact QR estimator $\hat{\theta}_{\tau,\text{QR}}$ as defined in equation (2), we have
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau,\text{QR}}-\theta_{\tau}\|$ $=O_{p}\left(\frac{1}{\sqrt{n}}\right),$ (12)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau,\text{QR}})\|_{p}$ $=O_{p}\left(\frac{k}{n}\right).$
@@ B.1 Auxiliary results
We use the maximal inequality in Ota et al. (2019) to establish (25). To do so, we verify the three conditions of this lemma. First,
@@
Consider $\hat{\theta}=\hat{\theta}_{\tau,p}$ defined in (3) for some $p\in[1,\infty]$ or $\hat{\theta}=\hat{\theta}_{\tau,\text{QR}}$, where $\tau\in(0,1)$. Under Assumptions 1, 2, and 3, $B_{n}(\hat{\theta})=B^{1}_{n}+B^{2}_{n}$, where the two components satisfy
@@ B.2 Proofs of main results on bias correction
$\hat{\kappa}_{\tau}$ $=\left(\tau-\frac{1}{2}\right)\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{n}\}}{2h_{n}}ZW^{\prime}G^{-1}Z\right]$
$+\sum_{j=1}^{k}\left(\tau-\frac{1}{2}\right)\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{n}\}}{2h_{n}}(e^{\prime}_{j}Z_{j})ZW^{\prime}\right](\hat{G}^{-1}-G^{-1})e_{j}$
$=\kappa_{\tau}+O_{p}\left(\frac{1}{\sqrt{n}}\right)+O_{p}\left(\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right)+o_{p}\left(\frac{\log h_{n}}{n^{5/6}h_{n}}\right)+O(h_{n}^{2})+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right).$
where the last term is based on the fact
$\|\hat{G}^{-1}-G^{-1}\|\leq\|\hat{G}^{-1}\|_{2}\|G^{-1}\|_{2}\|\hat{G}-G\|.$ (48)
So the overall rate is
$\max\left\{n^{-1/2},\,\frac{\sqrt{\log n}}{n^{2/5}},\,\frac{\log h_{n}}{n^{5/6}h_{n}},\,h_{n}^{2},\,\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right\}.$
The optimal bandwidth is $h_{3,n}\propto n^{-1/5}$ with
$\hat{\kappa}_{\tau}=\kappa_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right).$
@@ Appendix F Additional figures
In this section, we present additional simulation results. To explore the impact of the strength of the instruments, we consider two additional DGPs.
[…]
In DGP7, the instrument is weaker than in DGP1, and in DGP8, the instrument is stronger than in DGP4.
Figure 7: Bias (multiplied by $n$) before and after correction, IV strength grows from left to right
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for IVQR (implemented via the MILP formulation in Appendix D) without bias correction (blue dots), IVQR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and IVQR with infeasible bias correction (gold dashed line) for DGP7, DGP4, and DGP8 (ordered in terms of instrument strength). All results are based on 5,000 simulation repetitions. The infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and QR with infeasible bias correction (gold dashed line) for DGP1 with $n\in\{50,100,200\}$. All results are based on 5,000 simulation repetitions.
Figure 9: Bias (multiplied by $n$) before and after correction for DGP4, different sample sizes
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for IVQR (implemented via the MILP formulation in Appendix D) without bias correction (blue dots), IVQR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and IVQR with infeasible bias correction (gold dashed line) for DGP4 with $n\in\{50,100,200\}$. All results are based on 5,000 simulation repetitions. The infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
Figure 10: RMSE comparison of raw and bias-corrected estimators
Notes: The panels compare the RMSE for estimators without bias correction (blue), with infeasible bias correction (grey) and with feasible bias correction based on the rule-of-thumb bandwidth choice (gold) for (a) DGP2, classical QR and (b) DGP5, IVQR. All results are based on 5,000 simulation repetitions.
Figure 11: RMSE comparison of raw and bias-corrected estimators
Notes: The panels compare the RMSE for estimators without bias correction (blue), with infeasible bias correction (grey) and with feasible bias correction based on the rule-of-thumb bandwidth choice (gold) for (a) DGP3, classical QR and (b) DGP6, IVQR. All results are based on 5,000 simulation repetitions.
@@
Notes: The panels compare the Mean Absolute Deviation for the estimator without bias correction (blue) and with feasible bias correction (gold) for (a) DGP1, classical QR and (b) DGP4, IVQR. All results are based on 5,000 simulation repetitions.
Figure 13: Confidence interval coverage before and after correction for $n=100$
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP2 (classical QR) and DGP5 (IVQR). All results are based on 5,000 simulation repetitions.
Figure 14: Confidence interval coverage before and after correction for $n=100$
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP3 (classical QR) and DGP6 (IVQR). All results are based on 5,000 simulation repetitions.
Figure 15: Confidence interval coverage before and after correction for $n=200$
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP1 (classical QR) and DGP4 (IVQR). All results are based on 5,000 simulation repetitions.